OpenStudy (anonymous):

0 -6 This is an input and output table. Which 1 -7 rule describes how the output depends on X 2 -8 3 -9 A) -6x-1 B) -x^2-6 C) -x-6

OpenStudy (anonymous):

please help me. My class is almost over :/

OpenStudy (ivettef365):

Substitute x and find y to see if it is true

OpenStudy (ivettef365):

-x-6 -0-6 = -6 -1-6 = -7 -2-6 = -8 -3-6 = -9 then it is true